OpenStudy (anonymous):

Find the polynomial equation of least degree with roots -3, 0, and 3

OpenStudy (anonymous):

Is this Pre-Calculus?

OpenStudy (anonymous):

yes @KingThunder314

OpenStudy (solomonzelman):

The polynomial equation of least degree with roots \(\large\color{slate}{ \color{red}{a}, \color{green}{b}, \color{blue}{c} }\) would be written as: \(\large\color{slate}{ f(x)=(x-\color{red}{a})(x-\color{green}{b})(x-\color{blue}{c})}\)

OpenStudy (anonymous):

so the answer would be x^3+x^2+3x-9=0? @SolomonZelman

OpenStudy (solomonzelman):

For example, ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ roots \(\large\color{slate}{ \color{red}{4}, \color{green}{-3}, \color{blue}{5} }\) polynomial of least degree \(\large\color{slate}{ f(x)=(x-\color{red}{4})(x+\color{green}{3})(x-\color{blue}{5})}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ roots \(\large\color{slate}{ \color{red}{-2}, \color{green}{6}, \color{blue}{-10} }\) polynomial of least degree \(\large\color{slate}{ f(x)=(x+\color{red}{2})(x-\color{green}{6})(x+\color{blue}{10})}\)

OpenStudy (solomonzelman):

let me check

OpenStudy (solomonzelman):

nope, it is incorrect.

OpenStudy (anonymous):

Was I close?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

can you first write the f(x) in a similar to mine, form?

OpenStudy (anonymous):

f(x)=(x+3)(x-0)(x-3) @SolomonZelman

OpenStudy (solomonzelman):

yes that is right

OpenStudy (solomonzelman):

(x-0) is same as x, correct?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\large\color{slate}{ f(x)=x(x-3)(x+3) }\)

OpenStudy (anonymous):

Then FOIL?

OpenStudy (solomonzelman):

Lets focus on the \(\large\color{slate}{ (x-3)(x+3) }\) part for now. The difference of squares rule: \(\large\color{slate}{ (a-b)(a+b)=a^2-b^2 }\) Apply the difference of squares to the: \(\large\color{slate}{ (x-3)(x+3) }\) .

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x-3)(x+3) }\) becomes?

OpenStudy (anonymous):

x^3-9x=0 would be the right answer right?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

the function would be: \(\large\color{slate}{ f(x)=x^3-9x }\) just the polynomial is: \(\large\color{slate}{ x^3-9x }\) (it is all same, but in different form)

OpenStudy (solomonzelman):

You are very smart that you grabbed it....:)

OpenStudy (anonymous):

Thank you for the help! @SolomonZelman

OpenStudy (solomonzelman):

Sure, yw!